# Time-Varying Factor-Augmented Models for Volatility Forecasting

### v2

@@ Abstract
Keywords Volatility Forecasting, Factor Models, Long Short-Term Memory (LSTM), HAR, MIDAS, Volatility‑Based Pairs Trading
1 Introduction
From tick‑by‑tick hedging to long‑term investment strategy, reliable volatility forecasts are indispensable for guiding financial decisions. A vast literature—ranging from linear ARCH engle1982autoregressive , GARCH bollerslev1986generalized , and HAR corsi2009simple to nonlinear neural networks such as LSTMs hochreiter1997long —is typically applied on an asset-by-asset basis, modeling each asset in isolation. This univariate approach overlooks market-wide shocks that propagate rapidly across assets hamao1990correlations , often biasing forecasts and understating systemic risk. Although multivariate extensions such as BEKK‑GARCH engle1995multivariate , DCC‑GARCH engle2002dynamic , Conditional Autoregressive Wishart golosnoy2012conditional , and deep multivariate LSTM architectures capture shared dynamics, they are rarely used in real‑world settings due to parameter explosion, fragile estimation, and heavy tuning burdens bauwens2006multivariate ; engle2019large .
Factor models offer a natural remedy by efficiently summarizing co-movements through a low-dimensional set of latent drivers. Yet, two key limitations hinder their utility in volatility forecasting. First, most factor models are designed for returns, not volatility. Canonical linear frameworks such as Fama-French and Carhart fama1993common ; carhart1997persistence , and their modern machine‑learning successors gu2020empirical , all distill factors from returns data. Consequently, the rich patterns of commonality in volatility—which often exhibit their own distinct dynamics—remain largely unexplored. Second, although a few recent pioneering studies do extend factor models to realized volatility atak2013factor ; ding2025multiplicative , they still impose static factor loadings, which fail to capture time-varying interdependencies—an especially critical flaw given that volatility co-movement evolves over time del2008dynamic . Although recent econometric advances do provide time-varying loadings estimators mikkelsen2019consistent , these techniques have so far been confined exclusively to asset returns.
@@ 2 Related Work
Univariate Models. Frameworks from early AR models to more sophisticated ARCH engle1982autoregressive , GARCH bollerslev1986generalized ; bollerslev1987conditionally , and HAR corsi2009simple have long been pillars of the field due to their interpretability. Extensions have incorporated features like mixed-frequency data conrad2020two ; santos2014volatility and long-memory processes baillie1996long . However, by modeling each asset in isolation, these models suffer from a "univariate blind spot," ignoring the systemic co-movements that are crucial drivers of market-wide risk herskovic2016common ; hamao1990correlations .
Multivariate Models. Multivariate models are designed to capture asset interplay directly. Within the GARCH family, the BEKK specification engle1995multivariate guarantees positive-definite covariance matrices, while the Dynamic Conditional Correlation (DCC) model engle2002dynamic offers a more feasible estimation process. AR-style alternatives like the Conditional Autoregressive Wishart (CAW) golosnoy2012conditional provide further flexibility. While theoretically comprehensive, these models are often impractical for large systems due to the explosion of parameter count bauwens2006multivariate and numerically unstable estimation engle2019large .
Machine Learning Models. A parallel track has leveraged machine learning to capture the complex, nonlinear volatility patterns. Deep learning models like LSTM hochreiter1997long have shown superior performance over classical models by modeling long-range dependencies in financial time series liu2019novel ; mcnally2018predicting . However, these models are not immune to the "curse of dimensionality." When applied in a fully multivariate setting, they face excessive parameterization, high overfitting risk, and prohibitive computational costs ge2023review , inheriting the same fundamental limitations as their statistical counterparts.
Factor Models. Factor methods are the backbone of modern portfolio theory, distilling hundreds of correlated returns into a handful of common drivers and powering scalable asset allocation, hedging, and attribution. Widely used examples range from statistical factor‑GARCH variants such as OGARCH alexander2000primer and GO‑GARCH van2002go to recent factor‑structured deep networks feng2024deep . While these models achieve parsimony by using latent factors, most of the existing work targets returns gu2020empirical ; factor models for volatility remain comparatively sparse and always assume time‑invariant loadings atak2013factor ; ding2025multiplicative .
@@
3 Data and Volatility Measurement
3.1 Data
@@
3.2 Volatility Measurement
3.2.1 Daily Realized Volatility
Following andersen2001distribution , we estimate daily realized volatility (RV) nonparametrically from high-frequency midpoint prices. After filtering out non-positive, crossed, and spurious quotes, midpoint prices for asset $i$ on day $t$ are sampled every five minutes:
$\text{MID}_{i,t}=\frac{\text{BID}_{i,t}+\text{ASK}_{i,t}}{2}.$ (1)
@@ 3.2.1 Daily Realized Volatility
$RV_{i,t}=\sqrt{\sum_{s=1}^{N}\left[\ln\left(\frac{\text{MID}_{i,t_{s}}}{\text{MID}_{i,t_{s-1}}}\right)\right]^{2}}\;,$ (2)
@@
3.2.2 Aggregated Realized Volatility
To capture medium-term risk dynamics, we construct weekly realized volatility by averaging daily RVs over a 7-day window ($h=7$) following (corsi2009simple, ; bollerslev2018risk, ):
$RV_{t}^{7}=\frac{1}{7}\sum_{j=0}^{6}RV_{t-j}\;.$ (3)
@@ 3.2.2 Aggregated Realized Volatility
4 Methodology
4.1 Volatility Factor Analysis Across Assets
4.1.1 Time-Varying Factor Model
Let $\mathbf{y}_{t}=(RV_{1,t},\dots,RV_{p,t})$ denote the $p$-dimensional vector of realized volatilities at time $t=1,\ldots,T$. Following the time‑varying matrix factor model of chen2024timevaryingmatrixfactormodels , we posit the locally‑smooth factor representation:
$\mathbf{y}_{t}\;=\;\mathbf{\Lambda}_{t}\mathbf{f}_{t}+\mathbf{\varepsilon}_{t},$ (4)
@@ 4.1.1 Time-Varying Factor Model
$\mathbf{\Sigma}_{y,t}:=\mathbb{E}(\mathbf{y}_{t}\mathbf{y}_{t}^{\top})=\mathbf{\Lambda}_{t}\mathbf{\Sigma}_{f}\mathbf{\Lambda}_{t}^{\top}+\mathbf{\Sigma}_{\varepsilon}.$ (5)
Under mild regularity conditions, the rank‑$k$ signal component $\mathbf{\Lambda}_{t}\mathbf{\Sigma}_{f}\mathbf{\Lambda}_{t}^{\top}$ implies that the top $k$ eigenvectors of $\mathbf{\Sigma}_{y,t}$ span the column space of $\mathbf{\Lambda}_{t}$ up to rotation, as established in chen2024timevaryingmatrixfactormodels .
In principle, one could estimate $\mathbf{\Sigma}_{y,t}$ using kernel smoothing in $t/T$ as in chen2024timevaryingmatrixfactormodels . For practical implementation, we instead adopt a rolling‑window (uniform‑kernel) approximation. Fix a window length $n$ and define $W_{t}=\{t-n+1,\dots,t\}$. We estimate the local covariance matrix by
@@
4.1.2 Divergent Factor Structures: Concentration in Crypto vs. Diffusion in Equity
@@
In stark contrast, cryptocurrency’s risk is highly concentrated. A single, dominant systemic factor $\hat{\mathbf{f}}_{t}^{(1)}$ serves as the primary market-wide driver, accounting for a commanding 79.1% of total variance. This factor is closely linked to the Bitcoin cycle, with its overwhelming influence confirmed by the sharp "elbow" in Figure 1(d). While subsequent factors isolate the dynamics of major altcoins like Ethereum and Ripple, their marginal contributions are more modest, highlighting a market driven by one principal force.
These distinct factor structures provide a clear rationale for our horizon-dependent factor selection. For short-term (1-day) forecasts, we select only the single, dominant factor for each asset class to maximize the signal-to-noise ratio. For medium-term (7-day) forecasts, we employ an explained-variance threshold jolliffe2016principal ; pukthuanthong2009global . This involves selecting the minimum number of factors required to meet a predetermined cumulative variance target, which we tailor to each market’s unique structure (e.g., 85% for equities and 90% for crypto). This data-driven approach ensures we capture significant secondary dynamics without overfitting on noise.
@@
Figure 1: Factor Loadings and Scree Plots for Equities and Crypto
4.2 Factor-Augmented Forecasting Models
@@ 4.2.1 Autoregressive (AR)
$Y_{t+h}=\beta_{0}+\sum_{i=0}^{4}\beta_{i}RV_{t-i}+\epsilon_{t}.$ (6)
@@
$Y_{t+h}\;=\;\beta_{0}+\sum_{i=0}^{4}\beta_{i}RV_{t-i}\;+\;\sum_{j\in[S]}\gamma_{j}\,\widehat{\mathbf{f}}_{j,t}\;+\;\epsilon_{t}.$ (7)
4.2.2 Heterogeneous Autoregressive (HAR)
The HAR model captures multi-scale volatility persistence using daily, weekly, and monthly components corsi2009simple :
$Y_{t+h}=\beta_{0}+\beta^{d}RV_{t}+\beta^{w}RV^{7}_{t}+\beta^{m}RV^{30}_{t}+\epsilon_{t},$ (8)
@@ 4.2.2 Heterogeneous Autoregressive (HAR)
$\begin{split}Y_{t+h}\;=&\;\beta_{0}+\beta^{d}RV_{t}+\beta^{w}RV^{7}_{t}+\beta^{m}RV^{30}_{t}\\ &\;+\;\sum_{j\in[{S}^{d}]}\gamma^{d}_{j}\,\widehat{\mathbf{f}}^{\,d}_{j,t}\;+\;\sum_{r\in[{S}^{w}]}\gamma^{w}_{r}\,\widehat{\mathbf{f}}^{\,w}_{r,t}\;+\;\epsilon_{t},\end{split}$ (9)
@@
4.2.3 Mixed Data Sampling (MIDAS)
The MIDAS model captures long-memory dynamics by incorporating many past observations through a parsimonious weighting scheme, specified as a Beta-polynomial distributed lag ghysels2007midas :
$Y_{t+h}=\beta_{0}+\beta_{1}\left[a(1)^{-1}a(L)RV_{t}\right]+\epsilon_{t},$ (10)
@@ 4.2.3 Mixed Data Sampling (MIDAS)
$a_{i}=\left(\frac{i}{k}\right)^{\theta_{1}-1}\left(1-\frac{i}{k}\right)^{\theta_{2}-1}\frac{\Gamma(\theta_{1}+\theta_{2})}{\Gamma(\theta_{1})\Gamma(\theta_{2})},\quad i=1,\ldots,k,$ (11)
@@
$Y_{t+h}\;=\;\beta_{0}+\beta_{1}\!\bigl[a(1)^{-1}a(L)RV_{t}\bigr]+\sum_{j\in[S]}\gamma_{j}\bigl[b(1)^{-1}b(L)\,\widehat{\mathbf{f}}_{j,t}\bigr]+\epsilon_{t}.$ (12)
4.2.4 Long Short-Term Memory (LSTM)
@@ 4.3 Evaluation Metrics
4.3.1 Prediction Accuracy
Prediction accuracy is assessed with three widely used metrics: $R^{2}$, MSE, and QLIKE, which respectively capture explained variance, average prediction error, and distributional accuracy respectively patton2011volatility ; zhang2024volatility :
$R^{2}$ $=1-\sum_{t}\left(\frac{RV^{h}_{t+h}-\widehat{RV}^{h}_{t+h}}{RV^{h}_{t+h}-\overline{RV}^{h}_{t+h}}\right)^{2},$ (13)
$\mathrm{MSE}$ $=\frac{1}{T}\sum_{t}\left(RV^{h}_{t+h}-\widehat{RV}^{h}_{t+h}\right)^{2},$ (14)
$\mathrm{\mathrm{QLIKE}}$ $=\frac{1}{T}\sum_{t}\!\left[\frac{RV^{h}_{t+h}}{\widehat{RV}^{h}_{t+h}}-\ln\!\left(\frac{RV^{h}_{t+h}}{\widehat{RV}^{h}_{t+h}}\right)-1\right],$ (15)
4.3.2 Economic Implication

### v3

@@ Abstract.
Volatility Forecasting, Factor Models, Long Short-Term Memory (LSTM), HAR, MIDAS, Volatility‑Based Pairs Trading
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: 6th ACM International Conference on AI in Finance; November 15–18, 2025; Singapore, Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25), November 15–18, 2025, Singapore, Singaporedoi: 10.1145/3768292.3770407isbn: 979-8-4007-2220-2/2025/11ccs: Applied computing Operations researchccs: Computing methodologies Machine learning
1. Introduction
From tick‑by‑tick hedging to long‑term investment strategy, reliable volatility forecasts are indispensable for guiding financial decisions. A vast literature—ranging from linear ARCH (Engle, 1982), GARCH (Bollerslev, 1986), and HAR (Corsi, 2009) to nonlinear neural networks such as LSTMs (Hochreiter and Schmidhuber, 1997)—is typically applied on an asset-by-asset basis, modeling each asset in isolation. This univariate approach overlooks market-wide shocks that propagate rapidly across assets (Hamao et al., 1990), often biasing forecasts and understating systemic risk. Although multivariate extensions such as BEKK‑GARCH (Engle and Kroner, 1995), DCC‑GARCH (Engle, 2002), Conditional Autoregressive Wishart (Golosnoy et al., 2012), and deep multivariate LSTM architectures capture shared dynamics, they are rarely used in real‑world settings due to parameter explosion, fragile estimation, and heavy tuning burdens (Bauwens et al., 2006; Engle et al., 2019).
Factor models offer a natural remedy by efficiently summarizing co-movements through a low-dimensional set of latent drivers. Yet, two key limitations hinder their utility in volatility forecasting. First, most factor models are designed for returns, not volatility. Canonical linear frameworks such as Fama-French and Carhart (Fama and French, 1993; Carhart, 1997), and their modern machine‑learning successors (Gu et al., 2020), all distill factors from returns data. Consequently, the rich patterns of commonality in volatility—which often exhibit their own distinct dynamics—remain largely unexplored. Second, although a few recent pioneering studies do extend factor models to realized volatility (Atak and Kapetanios, 2013; Ding et al., 2025), they still impose static factor loadings, which fail to capture time-varying interdependencies—an especially critical flaw given that volatility co-movement evolves over time (Del Negro and Otrok, 2008). Although recent econometric advances do provide time-varying loadings estimators (Mikkelsen et al., 2019), these techniques have so far been confined exclusively to asset returns.
@@ 2. Related Work
Univariate Models. Frameworks from early AR models to more sophisticated ARCH (Engle, 1982), GARCH (Bollerslev, 1986, 1987), and HAR (Corsi, 2009) have long been pillars of the field due to their interpretability. Extensions have incorporated features like mixed-frequency data (Conrad and Kleen, 2020; Santos and Ziegelmann, 2014) and long-memory processes (Baillie, 1996). However, by modeling each asset in isolation, these models suffer from a ”univariate blind spot,” ignoring the systemic co-movements that are crucial drivers of market-wide risk (Herskovic et al., 2016; Hamao et al., 1990).
Multivariate Models. Multivariate models are designed to capture asset interplay directly. Within the GARCH family, the BEKK specification (Engle and Kroner, 1995) guarantees positive-definite covariance matrices, while the Dynamic Conditional Correlation (DCC) model (Engle, 2002) offers a more feasible estimation process. AR-style alternatives like the Conditional Autoregressive Wishart (CAW) (Golosnoy et al., 2012) provide further flexibility. While theoretically comprehensive, these models are often impractical for large systems due to the explosion of parameter count (Bauwens et al., 2006) and numerically unstable estimation (Engle et al., 2019).
Machine Learning Models. A parallel track has leveraged machine learning to capture the complex, nonlinear volatility patterns. Deep learning models like LSTM (Hochreiter and Schmidhuber, 1997) have shown superior performance over classical models by modeling long-range dependencies in financial time series (Liu, 2019; McNally et al., 2018). However, these models are not immune to the ”curse of dimensionality.” When applied in a fully multivariate setting, they face excessive parameterization, high overfitting risk, and prohibitive computational costs (Ge et al., 2023), inheriting the same fundamental limitations as their statistical counterparts.
Factor Models. Factor methods are the backbone of modern portfolio theory, distilling hundreds of correlated returns into a handful of common drivers and powering scalable asset allocation, hedging, and attribution. Widely used examples range from statistical factor‑GARCH variants such as OGARCH (Alexander, 2000) and GO‑GARCH (Van der Weide, 2002) to recent factor‑structured deep networks (Feng et al., 2024). While these models achieve parsimony by using latent factors, most of the existing work targets returns (Gu et al., 2020; CHEN et al., 2020; Chen et al., 2019; Chen and Chen, 2022; Liu and Chen, 2022; Chen and Fan, 2023; Chen et al., 2024d, a, c; Kong et al., 2024); factor models for volatility remain comparatively sparse and always assume time‑invariant loadings (Atak and Kapetanios, 2013; Ding et al., 2025).
@@
3. Data and Volatility Measurement
3.1. Data
@@
3.2. Volatility Measurement
3.2.1. Daily Realized Volatility
Following (Andersen et al., 2001), we estimate daily realized volatility (RV) nonparametrically from high-frequency midpoint prices. After filtering out non-positive, crossed, and spurious quotes, midpoint prices for asset $i$ on day $t$ are sampled every five minutes:
(1) $\text{MID}_{i,t}=\frac{\text{BID}_{i,t}+\text{ASK}_{i,t}}{2}.$
@@ 3.2.1. Daily Realized Volatility
(2) $RV_{i,t}=\sqrt{\sum_{s=1}^{N}\left[\ln\left(\frac{\text{MID}_{i,t_{s}}}{\text{MID}_{i,t_{s-1}}}\right)\right]^{2}}\;,$
@@
3.2.2. Aggregated Realized Volatility
To capture medium-term risk dynamics, we construct weekly realized volatility by averaging daily RVs over a 7-day window ($h=7$) following (Corsi, 2009; Bollerslev et al., 2018):
(3) $RV_{t}^{7}=\frac{1}{7}\sum_{j=0}^{6}RV_{t-j}\;.$
@@ 3.2.2. Aggregated Realized Volatility
4. Methodology
4.1. Volatility Factor Analysis Across Assets
4.1.1. Time-Varying Factor Model
Let $\mathbf{y}_{t}=(RV_{1,t},\dots,RV_{p,t})$ denote the $p$-dimensional vector of realized volatilities at time $t=1,\ldots,T$. Following the time‑varying matrix factor model of (Chen et al., 2024b), we posit the locally‑smooth factor representation:
(4) $\mathbf{y}_{t}\;=\;\mathbf{\Lambda}_{t}\mathbf{f}_{t}+\mathbf{\varepsilon}_{t},$
@@ 4.1.1. Time-Varying Factor Model
(5) $\mathbf{\Sigma}_{y,t}:=\mathbb{E}(\mathbf{y}_{t}\mathbf{y}_{t}^{\top})=\mathbf{\Lambda}_{t}\mathbf{\Sigma}_{f}\mathbf{\Lambda}_{t}^{\top}+\mathbf{\Sigma}_{\varepsilon}.$
Under mild regularity conditions, the rank‑$k$ signal component $\mathbf{\Lambda}_{t}\mathbf{\Sigma}_{f}\mathbf{\Lambda}_{t}^{\top}$ implies that the top $k$ eigenvectors of $\mathbf{\Sigma}_{y,t}$ span the column space of $\mathbf{\Lambda}_{t}$ up to rotation, as established in (Chen et al., 2024b).
In principle, one could estimate $\mathbf{\Sigma}_{y,t}$ using kernel smoothing in $t/T$ as in (Chen et al., 2024b). For practical implementation, we instead adopt a rolling‑window (uniform‑kernel) approximation. Fix a window length $n$ and define $W_{t}=\{t-n+1,\dots,t\}$. We estimate the local covariance matrix by
@@
4.1.2. Divergent Factor Structures: Concentration in Crypto vs. Diffusion in Equity
@@
In stark contrast, cryptocurrency’s risk is highly concentrated. A single, dominant systemic factor $\hat{\mathbf{f}}_{t}^{(1)}$ serves as the primary market-wide driver, accounting for a commanding 79.1% of total variance. This factor is closely linked to the Bitcoin cycle, with its overwhelming influence confirmed by the sharp ”elbow” in Figure 1(d). While subsequent factors isolate the dynamics of major altcoins like Ethereum and Ripple, their marginal contributions are more modest, highlighting a market driven by one principal force.
These distinct factor structures provide a clear rationale for our horizon-dependent factor selection. For short-term (1-day) forecasts, we select only the single, dominant factor for each asset class to maximize the signal-to-noise ratio. For medium-term (7-day) forecasts, we employ an explained-variance threshold (Jolliffe and Cadima, 2016; Pukthuanthong and Roll, 2009). This involves selecting the minimum number of factors required to meet a predetermined cumulative variance target, which we tailor to each market’s unique structure (e.g., 85% for equities and 90% for crypto). This data-driven approach ensures we capture significant secondary dynamics without overfitting on noise.
@@
Figure 1. Factor Loadings and Scree Plots for Equities and Crypto
4.2. Factor-Augmented Forecasting Models
@@ 4.2.1. Autoregressive (AR)
(6) $Y_{t+h}=\beta_{0}+\sum_{i=0}^{4}\beta_{i}RV_{t-i}+\epsilon_{t}.$
@@
(7) $Y_{t+h}\;=\;\beta_{0}+\sum_{i=0}^{4}\beta_{i}RV_{t-i}\;+\;\sum_{j\in[S]}\gamma_{j}\,\widehat{\mathbf{f}}_{j,t}\;+\;\epsilon_{t}.$
4.2.2. Heterogeneous Autoregressive (HAR)
The HAR model captures multi-scale volatility persistence using daily, weekly, and monthly components (Corsi, 2009):
(8) $Y_{t+h}=\beta_{0}+\beta^{d}RV_{t}+\beta^{w}RV^{7}_{t}+\beta^{m}RV^{30}_{t}+\epsilon_{t},$
@@ 4.2.2. Heterogeneous Autoregressive (HAR)
(9) $\begin{split}Y_{t+h}\;=&\;\beta_{0}+\beta^{d}RV_{t}+\beta^{w}RV^{7}_{t}+\beta^{m}RV^{30}_{t}\\ &\;+\;\sum_{j\in[{S}^{d}]}\gamma^{d}_{j}\,\widehat{\mathbf{f}}^{\,d}_{j,t}\;+\;\sum_{r\in[{S}^{w}]}\gamma^{w}_{r}\,\widehat{\mathbf{f}}^{\,w}_{r,t}\;+\;\epsilon_{t},\end{split}$
@@
4.2.3. Mixed Data Sampling (MIDAS)
The MIDAS model captures long-memory dynamics by incorporating many past observations through a parsimonious weighting scheme, specified as a Beta-polynomial distributed lag (Ghysels et al., 2007):
(10) $Y_{t+h}=\beta_{0}+\beta_{1}\left[a(1)^{-1}a(L)RV_{t}\right]+\epsilon_{t},$
@@ 4.2.3. Mixed Data Sampling (MIDAS)
(11) $a_{i}=\left(\frac{i}{k}\right)^{\theta_{1}-1}\left(1-\frac{i}{k}\right)^{\theta_{2}-1}\frac{\Gamma(\theta_{1}+\theta_{2})}{\Gamma(\theta_{1})\Gamma(\theta_{2})},\quad i=1,\ldots,k,$
@@
(12) $Y_{t+h}\;=\;\beta_{0}+\beta_{1}\!\bigl[a(1)^{-1}a(L)RV_{t}\bigr]+\sum_{j\in[S]}\gamma_{j}\bigl[b(1)^{-1}b(L)\,\widehat{\mathbf{f}}_{j,t}\bigr]+\epsilon_{t}.$
4.2.4. Long Short-Term Memory (LSTM)
@@ 4.3. Evaluation Metrics
4.3.1. Prediction Accuracy
Prediction accuracy is assessed with three widely used metrics: $R^{2}$, MSE, and QLIKE, which respectively capture explained variance, average prediction error, and distributional accuracy respectively (Patton, 2011; Zhang et al., 2024):
(13) $R^{2}$ $=1-\sum_{t}\left(\frac{RV^{h}_{t+h}-\widehat{RV}^{h}_{t+h}}{RV^{h}_{t+h}-\overline{RV}^{h}_{t+h}}\right)^{2},$
(14) $\mathrm{MSE}$ $=\frac{1}{T}\sum_{t}\left(RV^{h}_{t+h}-\widehat{RV}^{h}_{t+h}\right)^{2},$
(15) $\mathrm{\mathrm{QLIKE}}$ $=\frac{1}{T}\sum_{t}\!\left[\frac{RV^{h}_{t+h}}{\widehat{RV}^{h}_{t+h}}-\ln\!\left(\frac{RV^{h}_{t+h}}{\widehat{RV}^{h}_{t+h}}\right)-1\right],$
4.3.2. Economic Implication